OpenStudy (anonymous):

For the following function find the vertical, horizontal, and oblique asymptotes, if any. Please show all of your work. f(x) = 4x/7x^2+6x-16 +25. Can someone please help find the oblique asymptote?! For the following function find the vertical, horizontal, and oblique asymptotes, if any. Please show all of your work. f(x) = 4x/7x^2+6x-16 +25. Can someone please help find the oblique asymptote?! @Mathematics

OpenStudy (anonymous):

Hi, can you tell me what's confusing you, so we can figure out how to work on this together?

OpenStudy (anonymous):

Yes I dont know how to find the oblique asymptote

OpenStudy (anonymous):

You got it there \[6x+9\]

OpenStudy (anonymous):

Do you understand why?

OpenStudy (anonymous):

To find the oblique asymptote, you must use polynomial long division, and then analyze the function as it approaches infinity.

OpenStudy (anonymous):

It seems you used the division already. So you have to see what happens when x approach to infinity, and -infinity

OpenStudy (zarkon):

what is your function supposed to be \[f(x) = \frac{4x}{7x^2}+6x-16 +25\] \[f(x) = \frac{4x}{7x^2+6x-16 +25}\] \[f(x) = \frac{4x}{7x^2+6x-16 } +25\] or non of the above?

OpenStudy (zarkon):

*none

OpenStudy (anonymous):

Zarkon- the last one is what it is supposed to be

OpenStudy (anonymous):

This is what I have so far....now I need the interval notation for it....I suppose that is the oblique? to find the VA you must make the denominator = 0 7x^2+6x−16 factor to find the zeros (7x-8)(x+2)=0 VA are where this equation = 0 8/7, -2 to find the HA you have to find the limit as lim x→∞(4x÷(7x2+6x−16))+25 divide both top and bottom by the highest degree in the denominator which is x^2 simplify the expression now that we are left with 4/x in the numerator and 7+(6/x)-(16/x) in the denominator. Now we know that a constant over a infinitely larger variable will go to the y=0 axis. So this means that we are left with 0/7 +25. the equation of the HA is x=25. Because 0/7=0 and 0+25 =25.

OpenStudy (zarkon):

the equation is y=25 not x=25

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so what is the oblique asymptote?

OpenStudy (anonymous):

hello are you there?!